OpenStudy (anonymous):

Which equation is q=−2117 a solution for ? −17q=21 17q=21 21q=17 21q=−17

OpenStudy (anonymous):

none in my opinion unless you didnt type the whole problem

jimthompson5910 (jim_thompson5910):

I think you meant to say \[\Large q = -\frac{21}{17}\] ??

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

sorry about that.

jimthompson5910 (jim_thompson5910):

if you were to solve, say 21q=17, what do you get?

OpenStudy (anonymous):

q=17/21 ?

jimthompson5910 (jim_thompson5910):

so choice c isn't the answer

OpenStudy (anonymous):

oh okay

jimthompson5910 (jim_thompson5910):

you're looking for an equation where you get \[\Large q = -\frac{21}{17}\] as the solution